Solve by substitution.
6m+n=52
m-5n=19
Thank you,
Missy

Answer by Alan3354(69443)

Solve by substitution.
6m+n=52
m-5n=19 --> m = 5n+19
Sub for m in the 1st eqn
6m+n=52
6(5n+19) + n = 52
30n + 114 + n = 52
31n = -62
n = -2
---------
6m+n=52
6m - 2 = 52
6m = 54
m = 9